OpenStudy (anonymous):

How many solutions does the equation 2|x − 3| + 4 = 4 have? Zero One Two Infinitely Many

OpenStudy (anonymous):

That simplifies to abs(x-3)=0

OpenStudy (anonymous):

^ what does that mean @Luis_Rivera

OpenStudy (anonymous):

So theres only one solution right? @Xavier

OpenStudy (anonymous):

buttt 2|x − 3| + 4 = 4 given 2|x − 3| = 0 subtr 4 |x − 3| = 0 divide by 2 x=3 is the only solution

OpenStudy (anonymous):

No only one.

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

gee thanks. ^

OpenStudy (anonymous):

Your definition of the abolute value up there is wrong Luis abs(a)=a if a >= 0; -a if a < 0

OpenStudy (anonymous):

thanks lol. so wait it is only 1 solution right? cuz this was kinda hard to figure out.

OpenStudy (anonymous):

Thanks guys :)

OpenStudy (anonymous):

@Luis_Rivera GUESS WHAT BRO GUESS WHAT. I got my asnwers checked and.. YOU WERE WRONG. OHHH.